# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v2

@@ 7.5. Statistics
Component Lines of Lean
Pushout definition and lemmas 365
SVK theorem and infrastructure 1,004
Figure-eight application 330
2-sphere application 398
3-sphere (Hopf fibration) 393
Total (SVK-specific) 2,490
Note: This count covers only the SVK-specific modules. The full computational paths library—including path basics, rewrite system, homotopy theory, additional HITs (circle, torus, Klein bottle, projective plane, orientable surfaces), fibrations, Eilenberg-MacLane spaces, and the Hopf fibration—comprises approximately 25,100 lines.
7.6. Automation Potential
Since $\sim$ is decidable via normalization, a Lean tactic could automatically discharge $\pi_{1}$ equality goals by normalizing both sides and comparing. This would automate much of the tedious case analysis in proofs like Lemma 5.3. Future work could implement such a tactic using Lean 4’s metaprogramming facilities.
@@ 8. Conclusion
We have presented a formalization of the Seifert-van Kampen theorem within the computational paths framework, demonstrating that:
@@
    Pushouts can be defined as higher-inductive types with explicit path constructors.
@@
    The encode-decode method for computing $\pi_{1}(\mathrm{Pushout})$ works smoothly with computational paths, where path equality is captured by the decidable $\sim$ relation.
@@
    The SVK theorem yields practical applications: the figure-eight has non-abelian fundamental group $\mathbb{Z}\ast\mathbb{Z}$, and the 2-sphere is simply connected.
@@
    The entire development can be machine-checked in Lean 4, with axioms clearly identified.
@@ 8.1. Comparison with HoTT Approaches
Table 3. Comparison with HoTT Book SVK (§8.7)
@@
Path equality Identity type ($\mathrm{Id}$) Decidable ($\sim$)
Coherence witnesses Abstract existence Explicit rewrite derivations
Encode function Code family via $J$ Axiomatized
Decode function Direct construction Direct construction
$\pi_{1}$ computation Quotient Quotient by $\sim$
@@ 8.2. Limitations
    Axiom usage: The encode function and its properties are axiomatized. A full construction would require dependent elimination for HITs, not natively supported in Lean 4. In cubical Agda, these axioms would be provable.
@@
    Path-connectedness assumption: The SVK theorem requires path-connected types. For spaces with multiple components, a groupoid-valued version would be needed.
@@
    Rewrite system complexity: While the $\mathrm{LND}_{\mathrm{EQ}}$-TRS has 50+ rules, normalization complexity has not been formally analyzed.
@@
    Non-computational encode: Due to axiomatization, encode is not computationally evaluable. Decode, in contrast, is fully constructive.
@@
    $\sim$ triviality: As noted in Remark 2.5, our Lean implementation collapses $\sim$ via proof irrelevance, losing some intended distinctions between paths.
8.3. Recent Extensions
Since the initial development, the codebase has been extended with several features originally planned as future work:
@@ 8.3. Recent Extensions
    Higher homotopy groups $\pi_{n}$: Building on the weak $\omega$-groupoid structure, we now define:
    • Iterated loop spaces: $\texttt{Loop2Space}=\mathcal{D}_{2}(\rho_{a},\rho_{a})$, etc.
    • The second homotopy group $\pi_{2}(A,a)$ as the quotient of 2-loops by $\mathcal{D}_{3}$-connectivity
    • Eckmann-Hilton argument proving $\pi_{2}$ is abelian (piTwo_comm)
    • $\pi_{2}(S^{2})\cong 1$ via contractibility3
@@
    Truncation levels (n-types): Connecting the $\omega$-groupoid to HoTT truncation theory:
    • IsContr: contractible types (center + contraction paths)
    • IsProp: propositions (any two elements path-connected)
    • IsSet: sets (parallel paths are $\sim$)
    • IsGroupoid: 1-groupoids (parallel derivations connected)
    • Cumulative hierarchy: contractible $\to$ proposition $\to$ set $\to$ groupoid
    • All types are automatically 1-groupoids via contractibility3
@@
    Path simplification tactic: The path_simp tactic automates $\sim$ proofs:
    1macro "path_simp" : tactic =>
    2 ‘(tactic| simp only [rweq_refl, rweq_cmpA_refl_left,
    3 rweq_cmpA_refl_right, rweq_cmpA_inv_left, ...])
    Additional tactics include path_rfl, path_canon, and path_decide.
@@
    Covering space theory: Basic covering space infrastructure:
    • Total space $\Sigma_{a:A}P(a)$ for type families $P:A\to\mathcal{U}$
    • Path lifting via transport
    • $\pi_{1}(A,a)$ action on fibers via fiberAction
    • Universal cover with fiber $\pi_{1}(A,a)$
    • Deck transformations as covering automorphisms
  5. (5)
    Encode-decode improvements: The wedge encode-decode proofs now derive injectivity and surjectivity from the fundamental axioms, providing a cleaner interface.
  6. (6)
    Additional fundamental group computations: Beyond the figure-eight and 2-sphere:
    • Torus: $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$ via encode-decode
    • Real projective plane: $\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$
    • Orientable surfaces: $\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid\prod[a_{i},b_{i}]=1\rangle$
    • Möbius band and cylinder: $\pi_{1}\simeq\mathbb{Z}$ via deformation retract to circle
  7. (7)
    Fibrations and fiber sequences: Full fibration theory for computational paths:
    • Fiber types: $\mathrm{Fiber}(f,b):=\{a:A\mid f(a)=b\}$ with projection and inclusion
    • Fiber sequences $F\to E\to B$ with equivalence to fibers
    • Type families as fibrations with path lifting via transport
    • Connecting map $\partial:\pi_{1}(B,b)\to F$ via transport
    • Long exact sequence: $\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)\to\pi_{0}(F)$
    • Induced maps $f_{*}:\pi_{1}(A,a)\to\pi_{1}(B,f(a))$ with functoriality
    • Exactness proofs using $\Sigma$-type path characterization
  8. (8)
    Suspension-loop adjunction: Foundations for the adjunction $\Sigma X,Y]_{*}\cong[X,\Omega Y]_{*}$:
    • Pointed types and pointed maps with basepoint preservation
    • Suspension $\Sigma X$ as pointed type (north as basepoint)
    • Loop space $\Omega Y$ as pointed type (refl as basepoint)
    • Adjunction map: $f:\Sigma X\to Y$ yields $X\to\Omega Y$ via meridians
    • Basepoint preservation: $x_{0}\mapsto\rho$ when $f(\mathrm{north})=y_{0}$
    • Connectivity structures: IsPathConnectedPointed, IsSimplyConnected
  9. (9)
    Eilenberg-MacLane spaces $K(G,n)$: Characterization of classifying spaces:
    • Group and abelian group structures (GroupStr, AbelianGroupStr)
    • $K(G,1)$ characterization via IsKG1: path-connected, $\pi_{1}\cong G$, $\pi_{2}=1$
    • Circle is $K(\mathbb{Z},1)$: uses encode-decode with $\pi_{2}$ triviality from contractibility3
    • Loop space property: $\Omega(K(G,n+1))\simeq K(G,n)$ (stated)
    • Classifying spaces: $BG=K(G,1)$ structure
8.4. Remaining Future Work
  1. (1)
    Cubical type theory comparison: Relate computational paths to cubical approaches [1], exploring whether explicit rewrites complement cubical computation.
  2. (2)
    Fully constructive encode: Implement the code family to eliminate encode axioms, via cubical embedding or macro-generated eliminators.
  3. (3)
    Syntactic paths: Implement paths as syntactic terms (rather than equality-proof wrappers) to recover non-trivial $\sim$ distinctions.
  4. (4)
    Covering space classification: Prove the Galois correspondence between connected covers and subgroups of $\pi_{1}$.
  5. (5)
    Hopf fibration completion: The Hopf fibration framework is established with $\pi_{1}(S^{3})=1$ proved via SVK. Remaining work includes: constructing the Hopf map explicitly via the join $S^{1}*S^{1}\simeq S^{3}$; proving the equivalence $S^{3}\simeq\Sigma_{x:S^{2}}S^{1}$ without axioms; and extending to the quaternionic ($S^{3}\to S^{7}\to S^{4}$) and octonionic ($S^{7}\to S^{15}\to S^{8}$) Hopf fibrations.

### v3

@@ 8.7. Statistics
Metric Value
Total lines of Lean 41,130
Number of modules 107
Kernel axioms 36
Theorems/lemmas 1,200+
Definitions 500+
8.8. Assumption Manifest
For reviewer clarity, we provide a unified summary of all assumptions in the formalization.
Part A: Kernel Axioms (36 total, all for HIT type/constructor declarations)
HIT Kernel Axioms (Lean names)
Circle (7) Circle, circleBase, circleLoop, circleRec, circleRec_base, circleRec_loop, circleInd
Cylinder (11) Cylinder, cylinderBase0/1, cylinderSeg, cylinderLoop0/1, cylinderSurf, cylinderRec, cylinderRec_base0/1, cylinderRec_loop0
KleinBottle (5) KleinBottle, kleinBase, kleinLoopA/B, kleinSurf
OrientableSurface (5) OrientableSurface, base, loopA, loopB, surfaceRelation
NonOrientableSurface (4) NonOrientableSurface, nonOrientableBase, nonOrientableLoop, nonOrientableSurface2Cell
ProjectivePlane (4) ProjectivePlane, projectiveBase, projectiveLoop, projectiveLoopSquare
Not kernel axioms: Pushout (Quot-based), Sphere (suspension), Torus ($S^{1}\times S^{1}$), Wedge, Suspension, Lens spaces, Bouquet.
Part B: Typeclass Assumptions for SVK (Theorem 5.1)
Typeclass Role Status
HasGlueNaturalLoopRwEq Glue naturality Required; auto-derived for subsingletons
HasPushoutSVKEncodeQuot Encode map exists Assumed (user-provided)
HasPushoutSVKDecodeEncode Decode-encode round-trip Assumed (user-provided)
HasPushoutSVKEncodeDecode Encode-decode round-trip Assumed (user-provided)
HasPushoutSVKDecodeAmalgBijective Alternative: bijectivity only Assumed (choice-based)
Proved unconditionally: Decode map construction, decode respects amalgamation (given glue naturality).
Part C: Opt-in Axiom Instances (for assumption-free APIs)
These are additional kernel axioms that globally instantiate typeclasses:
  • instHasCirclePiOneEncodeAxiom in CirclePiOneAxiom.lean
  • instHasKleinPiOneEncodeAxiom in KleinPiOneAxiom.lean
  • instHasProjectivePiOneEncodeAxiom in ProjectivePiOneAxiom.lean
  • instHasLensPiOneEncodeAxiom in LensPiOneAxiom.lean
  • instHasWedgeSVKDecodeBijectiveAxiom in WedgeSVKAxiom.lean
These are not counted in the 36 core axioms; they are optional imports for convenience.
9. Library Highlights
Beyond the SVK framework, the library includes extended features for advanced homotopy-theoretic constructions:
9.1. Higher Homotopy Groups
The higher homotopy groups $\pi_{n}(A,a)$ for $n\geq 2$ are defined via iterated loop spaces:
$\pi_{n}(A,a):=\pi_{1}(\Omega^{n-1}(A,a),\rho)$
Key results include:
  • Abelianness of $\pi_{2}$ via Eckmann-Hilton (HigherHomotopy.lean)
  • Long exact sequence of a fibration (Fibration.lean)
9.2. Weak $\omega$-Groupoid Structure
Types are shown to form weak $\omega$-groupoids via the derivation tower:
  • 1-cells: paths $\mathrm{Path}(a,b)$
  • 2-cells: derivations $\mathcal{D}_{2}(p,q)$ witnessing $p\sim q$
  • 3-cells: derivations $\mathcal{D}_{3}(d_{1},d_{2})$ between derivations
  • And so on, with truncation at level 3 (OmegaGroupoid.lean)
9.3. Truncation Levels
The library formalizes HoTT-style truncation levels:
  • IsContr: Contractible types (all points equal)
  • IsProp: Propositions (all proofs equal)
  • IsSet: Sets (all paths equal)
  • IsGroupoid: Groupoids (all 2-paths equal)
Module: Truncation.lean
9.4. Covering Spaces
Covering space theory with $\pi_{1}$-actions:
  • Fiber transport action (CoveringSpace.lean)
  • Galois correspondence framework (CoveringClassification.lean)
9.5. Additional Features
  • Eilenberg-MacLane spaces: $S^{1}=K(\mathbb{Z},1)$ characterization (EilenbergMacLane.lean)
  • Mayer-Vietoris: Abelianized SVK sequence (MayerVietoris.lean)
  • Path tactics: Automation via path_simp, path_auto, path_normalize (PathTactic.lean)
@@ 10. Conclusion
We have presented a modular SVK framework within the computational paths setting, demonstrating that:
@@
    Pushouts can be implemented via Lean’s quotient types with zero kernel axioms, while computation rules are supplied as typeclass assumptions.
@@
    The encode-decode method works smoothly with split assumptions, supporting both data-level and Prop-level (choice-based) interfaces. The SVK equivalence is parametric in user-supplied encode/decode structure.
@@
    The framework yields instantiations for all closed surfaces, lens spaces, projective spaces, and more—with explicit dependency tracking via typeclasses.
@@
    The entire development comprises 41,130 lines of Lean 4 code with only 36 kernel axioms (for HIT type/constructor declarations).
@@ 10.1. Comparison with HoTT Approaches
Table 3. Comparison with Standard HoTT
@@
Path equality Identity type ($\mathrm{Id}$) Syntactic ($\sim$)
Coherence Abstract existence Explicit rewrite derivations
SVK assumptions Monolithic Split/modular typeclasses
Encode function Code family via $J$ Typeclass assumption
$\omega$-groupoid Higher identity types Derivation tower
Decidable under termination + confluence; termination not mechanized here (see §2.4).
@@ 10.2. Limitations
    Axiom usage: The encode/decode equivalences are axiomatized via typeclasses rather than proved. In cubical Agda, where transport computes, these would be provable. This is an inherent limitation of standard Lean’s proof-irrelevant Prop—we cannot prove that transport along a non-trivial loop acts non-trivially. See Remark 8.4 for why this design is nonetheless consistent.
@@
    Explicit step storage: Paths store explicit List (Step A) to distinguish loops—a departure from the “thin” identity types of standard MLTT. This is essential for non-trivial $\pi_{1}$ (see Remark 2.1).
@@
    Path-connectedness assumption: The SVK theorem requires path-connected types.
10.3. Future Work
@@
    Cubical type theory comparison: Relate computational paths to cubical approaches.
@@
    Fully constructive encode: Implement code families to eliminate encode axioms.
@@
    Covering space classification: Prove the Galois correspondence between covering spaces and subgroups of $\pi_{1}$.
@@
    Higher Hopf fibrations: Quaternionic ($S^{3}\to S^{7}\to S^{4}$) and octonionic.
@@
    3-manifold classification: Seifert fibered spaces, graph manifolds.
@@
    Homology theory: Extend from $\pi_{1}$ to full homology groups.
Appendix A Application Inventory
The following table provides the exact Lean theorem names and module paths for the headline results. All theorems depend on typeclass assumptions as indicated; “opt-in” means an axiom import is available for assumption-free use.
Result Lean Name Module Deps
$\pi_{1}(S^{1})\simeq\mathbb{Z}$ circlePiOneEquivInt CircleStep Opt-in
$\pi_{1}(T^{2})\simeq\mathbb{Z}^{2}$ torusPiOneEquivIntProd TorusStep Opt-in
$\pi_{1}(S^{2})\simeq 1$ sphere2PiOneEquivUnit Sphere None
$\pi_{1}(S^{1}\vee S^{1})$ figureEightPiOneEquiv FigureEight SVK
SVK schema seifertVanKampenEquiv PushoutPaths SVK
SVK choice svkEquiv_of_bijective PushoutPaths Bij.
$\pi_{1}(K)$ kleinPiOneEquivSemidirect KleinBottleStep Opt-in
$\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$ projectivePiOneEquivZ2 ProjectivePlane Opt-in
$\pi_{1}(\Sigma_{g})$ piOneEquivPresentation OrientableSurface SVK
$\pi_{1}(N_{n})$ piOneEquivPresentation NonOrientableSurface SVK
$\pi_{1}(L(p,q))$ lensPiOneEquivZMod LensSpace Opt-in
Wedge wedgeFundamentalGroupEquiv WedgePaths Opt-in
Legend: Opt-in: axiom import provides global instance. SVK: requires encode/decode typeclasses. Bij.: requires HasPushoutSVKDecodeAmalgBijective. None: no extra assumptions.